OpenStudy (anonymous):

evaluate 5^28.5^-32.5^-4

OpenStudy (anonymous):

5^-8

OpenStudy (turingtest):

please use parentheses to clarify the question

OpenStudy (anonymous):

Its either 17,920 B 0.017920 C 0.000256 D 0.00000256

OpenStudy (turingtest):

is it (5^28)(5^-32)(5^-4)? then we have 5^-8=2.56x10^-6 which is D

jhonyy9 (jhonyy9):

us the exponents property when divide the same terms with exponents you need assuming those exponents - so in this case there we have 28-32-4=28-36=-8 that like resulte we have got 5 on exponent -8